# Block-Diagonal Orthogonal Relation and Matrix Entity for Knowledge Graph Embedding

### v2

@@ 4.4 Optimization
Traditional KGE models train and optimize relations and entities jointly. In contrast, our study aims for more effective optimization of the relation embeddings’ block-diagonal orthogonal matrices $\mathbf{X}_{i}\in\mathbb{R}^{d\times d}$ by separately optimizing relations and entities, utilizing Riemannian optimization for the relation matrix $\mathbf{e}_{R}\in\mathbb{R}^{n\times n}$ and SGD for the entity matrix $\mathbf{e}_{V}\in\mathbb{R}^{n\times m}$.
Initially, when optimizing relations, all entity parameters are fixed, rendering the entity embeddings analogous to the function $f(\cdot)$ in the problem 1. Notably, each block-diagonal orthogonal matrix $\mathbf{X}_{i}$ within the relation embedding $\mathbf{e}_{R}$ optimized by individual Riemannian optimization using RiemannianAdam Kochurov et al. (2020), which is a Riemannian version (equation 2) of the popular Adam optimizer Kingma and Ba (2014). These are then concatenated in a block-diagonal way according to equation 3 to complete the process. After optimizing the relation parameters $\mathbf{e}_{R}\in\mathbb{R}^{n\times n}$, they are held constant while the entity parameters $\mathbf{e}_{V}\in\mathbb{R}^{n\times m}$ are optimized using Stochastic Gradient Descent (SGD), specifically employing the Adagrad optimizer Duchi et al. (2011).
@@ 5 Experiment
We expect that our proposed OrthogonalE model, employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations, will outperform baseline models. Also, we anticipate OrthogonalE is a general and flexible KGE model and it can represent several relation patterns simultaneously. Our goal is to validate these through empirical testing.
@@ 5.1 Experiment Setup
We evaluate our proposed method on two KG datasets, including WN18RR Dettmers et al. (2018) (license: Apache 2.0), FB15K-237 Toutanova and Chen (2015) (license: CC-BY-4.0). The details of these datasets are shown in Table 4. More detail refers to A.1.
@@
To predict the tail entity from a given head entity and relation, we rank the correct tail entity among all possible entities using two established ranking metrics. First is the Mean Reciprocal Rank (MRR), the average inverse ranking of correct entities, calculated as $\frac{1}{n}\sum_{i=1}^{n}\frac{1}{\text{ Rank }_{i}}$. Second is Hits@K for $K\in\{1,3,10\}$, the frequency of correct entities ranking within the top $K$ positions.
@@
We compare our new model with classic methods, including TransE Bordes et al. (2013), DistMult Yang et al. (2014), ConvE Dettmers et al. (2018), and rotation-based KGE methods such as ComplEx Trouillon et al. (2016), RotatE Sun et al. (2019), QuatE Zhang et al. (2019), HopfE Bastos et al. (2021), and DensE Lu et al. (2022). In addition to these methods and our OrthogonalE$(d\times d)$, we introduce comparative models Gram-Schmidt$(d\times d)$ utilizing the Gram-Schmidt process for generating orthogonal matrices and SGD for joint relation-entity training. OrthogonalE further differentiates by employing orthogonal matrices of varying sizes to discuss performance nuances.
@@
The key hyperparameters in our implementation include the learning rate for RiemannianAdam Kochurov et al. (2020) and Adagrad Duchi et al. (2011), negative sample size, and batch size. To determine the optimal hyperparameters, we performed a grid search using the validation data. More detail refers to A.1.
Figure 4: MRR accuracy comparison of OrthogonalE models with different small orthogonal matrices across varying entity dimensions ($n*1$, we set $m=1$) on WN18RR and FB15K-237.
@@ 5.2.1 Overall Accuracy
5.2.2 Block-diagonal Orthogonal matrices
Fig. 4 shows MRR accuracy comparison of OrthogonalE models with different block-diagonal orthogonal matrices across varying entity dimensions ($n*1$, we set $m=1$) on WN18RR and FB15K-237.
An initial dataset analysis reveals WN18RR has 40,943 entities with just 11 relations (about 3,722 entities per relation), while FB15K-237 includes 14,541 entities and 237 relations (around 61 entities per relation). This implies WN18RR requires a more sophisticated representation capability compared to FB15K-237.
Our results (Fig. 4) confirm our dataset analysis. For WN18RR, performance is similar for block sizes from 3$\times$3 to 10$\times$10, all outperforming 2$\times$2 blocks, showcasing 2$\times$2 blocks are not enough for its relation representation. However, for FB15K-237, performance is stable across all block sizes, indicating 2$\times$2 blocks are enough for its relations representation. These results show WN18RR requires more complex blocks for adequate representation, and illustrate that the OrthogonalE model is general, which can adapt to datasets of various complexities by adjusting the dimension $d$ of the block-diagonal matrices.
@@ 5.2.2 Block-diagonal Orthogonal matrices
Figure 6: MRR accuracy comparison of RotatE and OrthogonalE(2$\times$2) models across varying entity dimensions ($n*m$) on WN18RR.
Figure 7: MRR accuracy comparison of QuatE and OrthogonalE(3$\times$3) models across varying entity dimensions ($n*m$) on WN18RR.
Figure 8: MRR accuracy comparison of RotatE and OrthogonalE(2$\times$2) models across varying entity dimensions ($n*m$) on FB15K-237.
@@ 5.2.3 Riemannian Optimization for relations
Fig. 5 compares MRR accuracies of OrthogonalE (2$\times$2) and Gram-Schmidt (2$\times$2) across entity dimensions ($m$) with a constant relation matrix (40$\times$40) on WN18RR, assessing the efficacy of orthogonal optimization beyond the Gram-Schmidt method for block-diagonal orthogonal matrices. The result demonstrates that OrthogonalE’s Riemannian optimization significantly exceeds Gram-Schmidt, underscoring its necessity.
5.2.4 Entity matrix
In OrthogonalE, we maintained a constant entity dimension ($n*m$) while varying $m$ to assess the impact of entity shape. Fig. 6 compare MRR accuracies of RotatE with OrthogonalE (2$\times$2) over different fixed entity dimensions $n*m$ on WN18RR. OrthogonalE models with $m=1$, $2$, or $3$ perform similarly and better than $m=10$, and all significantly outperform RotatE across dimensions. Notably, their relation parameter is $1/m$ of RotatE’s, which shown in Table. 2. These results demonstrate OrthogonalE’s efficacy in saving relation parameters while outperforming RotatE, highlighting our model’s flexibility in controlling entity dimension through variable $m$ without unnecessarily increasing relation size.
Besides the comparison of RotatE and OrthogonalE(2$\times$2), Fig. 7 compare MRR accuracies of QuatE and OrthogonalE (3$\times$3) over different fixed entity dimensions $n*m$ on WN18RR, then Fig. 8 shows comparison of RotatE and OrthogonalE (2$\times$2) in FB15K-237. The experimental results, consistent with those discussed in the previous paragraph, demonstrate our model’s flexibility in controlling entity dimension through variable $m$ without unnecessarily increasing relation size.
@@ 5.2.5 Relation Pattern
Following the proof of relation patterns in Appendix A.2, Fig. 9 shows histograms of relation embeddings for different relation patterns. We provide several relation patterns examples and discussion of non-commutative composition in Appendix A.3
@@
In OrthogonalE, the symmetry relation pattern is encoded when the $\mathbf{e}_{R}$ embedding satisfies $\mathbf{e}_{R}\cdot\mathbf{e}_{R}=\mathrm{I}$, in accordance with Equation LABEL:equ:symmetery_proof. Figs. 9(a) and (b) illustrate the embeddings of $\mathbf{e}_{R}^{1}$ and $\mathbf{e}_{R}^{1}\cdot\mathbf{e}_{R}^{1}-\mathrm{I}$, respectively. From Fig. 9(b), we observe that nearly all values are concentrated around 0, thereby indicating that OrthogonalE’s relations exhibit symmetry properties. Correspondingly, the multitude of non-zero values in Fig. 9(d) indicates that OrthogonalE’s relations also can represent antisymmetry properties.
@@
The inversion relation pattern is encoded when the $\mathbf{e}_{R}^{3}$ and $\mathbf{e}_{R}^{4}$ embedding satisfy $\mathbf{e}_{R}^{3}\cdot\mathbf{e}_{R}^{4}=\mathrm{I}$, in accordance with Equation LABEL:equ:inversion_proof. Even though $\mathbf{e}_{R}^{3}$ and $\mathbf{e}_{R}^{4}$ are responsible for additional relation patterns, which results in a cluster of values around -2 in Fig 9 (g), the majority of values still converge towards or equal 0. This suggests that OrthogonalE’s relations have the inversion property.
@@
The composition relation pattern is encoded when the $\mathbf{e}_{R}^{5}$, $\mathbf{e}_{R}^{6}$, and $\mathbf{e}_{R}^{7}$ embedding satisfy $\mathbf{e}_{R}^{6}\cdot\mathbf{e}_{R}^{5}=\mathbf{e}_{R}^{7}$, in accordance with Equation LABEL:equ:composition_proof. The majority of data in Fig. 9 (k) converge towards or are equal to 0, indicating that OrthogonalE’s relations can represent the composition relation pattern.
@@ 6 Conclusion
In this study, we propose the OrthogonalE model to acquire a flexible and general KGE model with employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations. Experimental results indicate that our new KGE model, OrthogonalE, offers generality and flexibility, captures several relation patterns, and significantly outperforms state-of-the-art KGE models while substantially reducing the number of relation parameters.
@@
Even though the Block-Diagonal Orthogonal relation with Riemannian Optimization makes KGE models more general and improves their performance, the computation of Exponential Retraction in the Orthogonal manifold for Riemannian Optimization is very costly. In practical model training, with the same entity dimension, our OrthogonalE (2$\times$2) training time is 4 times longer than that of RotatE. In future research directions, we will continue to explore this limitation, such as by employing the landing algorithm Ablin and Peyré (2022) for retraction on orthogonal manifolds to reduce computational complexity.
@@ A.1 More information about Experiment setup
WN18RR is a subset of WN18 Dettmers et al. (2018) which is contained in WordNet Miller (1995). FB15K-237 is a subset of FB15K, whichh is a subset of Freebase Bollacker et al. (2008), a comprehensive KG including data about common knowledge. All three datasets were designed for KGE, and we employ them for KGE tasks, and all three datasets have no individual people or offensive content.
@@ A.4 Introduction of Relation Patterns
Certain relations demonstrate symmetry, indicating that a relation’s validity between entities $x$ and $y$ ($\left(r_{1}(x,y)\Rightarrow r_{1}(y,x)\right)$) (for instance, is married to) is equally valid in the opposite direction (namely, from $y$ to $x$). Conversely, other relations are characterized by antisymmetry ($\left(r_{1}(x,y)\Rightarrow\neg r_{1}(y,x)\right)$), signifying that if a relation is applicable between $x$ and $y$ (such as is father of), it is inapplicable in the reverse direction (from $y$ to $x$).
@@
HittERChen et al. (2020) Language Models .503
@@ A.5 Other baseline KGE model
In recent times, several significant performance methods have been developed, as detailed for WN18RR in Table 6. Among these, MoCoSAHe et al. (2023), SimKGCWang et al. (2022a), C-LMKEWang et al. (2022b), KNN-KGEZhang et al. (2022), and HittERChen et al. (2020) mainly utilize Language Models (LMs) to enrich dataset semantic information, thereby achieving superior outcomes. Conversely, LERPHan et al. (2023) does not employ LMs but leverages additional contextual information (logic rules) beyond the dataset to fill information gaps in entities and relations. On the other hand, methods such as TransEBordes et al. (2013), RotatESun et al. (2019), and the OrthogonalE method introduced in this paper depend solely on the inherent data and information of the KGE dataset itself. These methods, based on specific mathematical rules and algorithms, do not incorporate any external information and thus do not operate as black-box approaches like LLMs. Consequently, these dataset-dependent methods remain highly valuable for KGE research.
@@
All the hyperparameter settings have been shown in Table 5

### v3

@@ 4.4 Optimization
Traditional KGE models train and optimize relations and entities jointly. In contrast, our study aims to achieve more effective optimization of the block-diagonal orthogonal matrices of relation embeddings $\mathbf{X}_{i}\in\mathbb{R}^{d\times d}$ by separately optimizing relations and entities, utilizing Riemannian optimization for the relation matrix $\mathbf{e}_{R}\in\mathbb{R}^{n\times n}$ and SGD for the entity matrix $\mathbf{e}_{V}\in\mathbb{R}^{n\times m}$.
Initially, when optimizing relations, all entity parameters are fixed, rendering the entity embeddings analogous to the function $f(\cdot)$ in the problem formulated in (1). Notably, each block-diagonal orthogonal matrix $\mathbf{X}_{i}$ within the relation embedding $\mathbf{e}_{R}$ optimized by individual Riemannian optimization of eq. (2) using RiemannianAdam Kochurov et al. (2020), which is a Riemannian version of the popular Adam optimizer Kingma and Ba (2014). These are then concatenated in a block-diagonal way according to eq. (3) to complete the process. After optimizing the relation parameters $\mathbf{e}_{R}\in\mathbb{R}^{n\times n}$, they are held constant while the entity parameters $\mathbf{e}_{V}\in\mathbb{R}^{n\times m}$ are optimized using Stochastic Gradient Descent (SGD), specifically employing the Adagrad optimizer Duchi et al. (2011).
@@ 5 Experiment
We expect that our proposed OrthogonalE model, employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations, will outperform baseline models. Also, we anticipate that OrthogonalE is a general and flexible KGE model and can represent several relation patterns simultaneously. Our goal is to validate these through empirical testing.
@@ 5.1 Experiment Setup
We evaluate our proposed method on two KG datasets, including WN18RR Dettmers et al. (2018) (license: Apache 2.0), FB15K-237 Toutanova and Chen (2015) (license: CC-BY-4.0). The details of these datasets are shown in Table 4. More detail is given in Appendix A.1.
@@
To predict the tail entity from a given head entity and relation, we rank the correct tail entity among all possible entities using two established ranking metrics. The first is the mean reciprocal rank (MRR), the average inverse ranking of the correct entities, calculated as $\frac{1}{n}\sum_{i=1}^{n}\frac{1}{\text{ Rank }_{i}}$. The second is Hits@$K$ for $K\in\{1,3,10\}$, the frequency of correct entities ranking within the top $K$ positions.
@@
We compare our new model with several classic methods, including TransE Bordes et al. (2013), DistMult Yang et al. (2014), ComplEx Trouillon et al. (2016), and ConvE Dettmers et al. (2018). Additionally, we include rotation-based KGE methods such as RotatE Sun et al. (2019), QuatE Zhang et al. (2019), HopfE Bastos et al. (2021), and DensE Lu et al. (2022) as baselines. In addition to these methods and our OrthogonalE$(d\times d)$, we introduce comparative models such as Gram-Schmidt$(d\times d)$ utilizing the Gram-Schmidt process for generating orthogonal matrices and SGD for joint relation-entity training. OrthogonalE further differentiates by employing orthogonal matrices of varying sizes to discuss performance nuances.
@@
The key hyperparameters of our implementation include the learning rate for RiemannianAdam Kochurov et al. (2020) and Adagrad Duchi et al. (2011), negative sample size, and batch size. To determine the optimal hyperparameters, we performed a grid search using the validation data. More detail refers to Appendix A.1.
Figure 4: MRR accuracy comparison of OrthogonalE models with different block-diagonal orthogonal matrices across varying entity dimensions ($n\times 1$, where we set $m=1$ to control the entity shape as a single entity vector) on WN18RR and FB15K-237.
@@ 5.2.1 Overall Accuracy
5.2.2 Block-diagonal Orthogonal matrices for generality
To prove the generality of OrthogonalE, we design experiments to compare the MRR accuracy of OrthogonalE models with different block-diagonal orthogonal matrices across varying entity dimensions ($n\times 1$, where we set $m=1$ to control the entity shape as a single entity vector) on WN18RR and FB15K-237. The results are shown in Fig. 4 and the explanation is provided as follows.
An initial dataset analysis reveals WN18RR has 40,943 entities with just 11 relations (about 3,722 entities per relation), while FB15K-237 includes 14,541 entities and 237 relations (around 61 entities per relation). This implies that WN18RR requires a more sophisticated representation capability compared to FB15K-237.
Our results (Fig. 4) confirm our dataset analysis. For WN18RR, the performance is similar for block sizes from 3$\times$3 to 10$\times$10, all outperforming 2$\times$2 blocks, showcasing 2$\times$2 blocks are not enough for its relation representation. However, for FB15K-237, performance is stable across all block sizes, indicating 2$\times$2 blocks are enough for its relations representation. These results show WN18RR requires more complex blocks for adequate representation, and illustrate that the OrthogonalE model is general, which can adapt to datasets of various complexities by adjusting the dimension $d$ of the block-diagonal matrices.
@@ 5.2.2 Block-diagonal Orthogonal matrices for generality
Figure 6: MRR accuracy comparison of RotatE and OrthogonalE(2$\times$2) models across varying entity dimensions ($n\times m$) on WN18RR.
Figure 7: MRR accuracy comparison of RotatE and OrthogonalE(2$\times$2) models across varying entity dimensions ($n\times m$) on FB15K-237.
Figure 8: MRR accuracy comparison of QuatE and OrthogonalE(3$\times$3) models across varying entity dimensions ($n\times m$) on WN18RR.
@@ 5.2.3 Riemannian Optimization for relations
Fig. 5 compares MRR accuracies of OrthogonalE (2$\times$2) and Gram-Schmidt (2$\times$2) across entity dimensions ($m$) with a constant relation matrix (40$\times$40) on WN18RR, assessing the efficacy of orthogonal optimization beyond the Gram-Schmidt method for block-diagonal orthogonal matrices. The result demonstrates that OrthogonalE’s Riemannian optimization significantly exceeds Gram-Schmidt, underscoring its necessity. In addition, OrthogonalE can solve the problem that Gram-Schmidt model will face the gradient vanishing during the training process.
5.2.4 Entity matrix for flexibility
In OrthogonalE, we maintained a constant entity dimension ($n\times m$) while varying $m$ to assess the impact of entity shape. To prove the flexibility of OrthogonalE, we design experiments to compare the MRR accuracies of RotatE with OrthogonalE (2$\times$2) and QuatE and OrthogonalE (3$\times$3) over different fixed entity dimensions $n\times m$ in WN18RR and FB15K-237 dataset.
From the results on WN18RR in Fig. 6, OrthogonalE models with $m=1$, $2$, or $3$ perform similarly and better than $m=10$, and all significantly outperform RotatE across dimensions. Notably, their relation parameter count is $1/m$ of RotatE’s, which is shown in Table. 2. These results demonstrate OrthogonalE’s efficacy in saving relation parameters while outperforming RotatE, highlighting our model’s flexibility in controlling entity dimension through variable $m$ without unnecessarily increasing relation size.
Furthermore, Fig. 7 shows comparison of RotatE and OrthogonalE (2$\times$2) in FB15K-237. Fig. 8 compare MRR accuracies of QuatE and OrthogonalE (3$\times$3) on WN18RR. The experimental results, consistent with those discussed in the previous paragraph.
@@ 5.2.5 Relation Pattern
Following the proof of relation patterns in Appendix A.2, Fig. 9 shows histograms of relation embeddings for different relation patterns. We provide several examples of relation patterns and a discussion of non-commutative composition in Appendix A.3.
@@
In OrthogonalE, the symmetry relation pattern is encoded when the $\mathbf{e}_{R}$ embedding satisfies $\mathbf{e}_{R}\cdot\mathbf{e}_{R}=\mathrm{I}$, in accordance with eq. (LABEL:equ:symmetery_proof). Figs. 9(a) and (b) illustrate the embeddings of $\mathbf{e}_{R}^{1}$ and $\mathbf{e}_{R}^{1}\cdot\mathbf{e}_{R}^{1}-\mathrm{I}$, respectively. From Fig. 9(b), we observe that nearly all values are concentrated around 0, thereby indicating that OrthogonalE’s relations exhibit symmetry properties. Correspondingly, the multitude of nonzero values in Fig. 9(d) indicates that OrthogonalE’s relations also can represent antisymmetry properties.
@@
The inversion relation pattern is encoded when the $\mathbf{e}_{R}^{3}$ and $\mathbf{e}_{R}^{4}$ satisfies $\mathbf{e}_{R}^{3}\cdot\mathbf{e}_{R}^{4}=\mathrm{I}$, according to eq. (LABEL:equ:inversion_proof). Even though $\mathbf{e}_{R}^{3}$ and $\mathbf{e}_{R}^{4}$ are responsible for additional relation patterns, which results in a cluster of values around $-2$ in Fig 9 (g), the majority of values still converge towards or equal 0. This suggests that OrthogonalE’s relations have the inversion property.
@@
The composition relation pattern is encoded when the $\mathbf{e}_{R}^{5}$, $\mathbf{e}_{R}^{6}$, and $\mathbf{e}_{R}^{7}$ embedding satisfy $\mathbf{e}_{R}^{6}\cdot\mathbf{e}_{R}^{5}=\mathbf{e}_{R}^{7}$, in accordance with eq. (LABEL:equ:composition_proof). The majority of data in Fig. 9 (k) converge towards or are equal to 0, indicating that OrthogonalE’s relations can represent the composition relation pattern.
@@ 6 Conclusion
In this study, we propose the OrthogonalE model to acquire a flexible and general KGE model with employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations. Experimental results indicate that our new KGE model offers generality and flexibility, captures several relation patterns, and outperforms SoTA rotation-based KGE models while substantially reducing the number of relation parameters.
@@
Even though the block-diagonal orthogonal relation with Riemannian optimization makes KGE models more general and improves their performance, the computation of exponential retraction in the orthogonal manifold for Riemannian optimization is costly. In practical model training, with the same entity dimension, our OrthogonalE (2$\times$2) training time is 4 times longer than that of RotatE. In future research directions, we will continue to explore this limitation, such as by employing the landing algorithm Ablin and Peyré (2022) for retraction on orthogonal manifolds to reduce computational complexity.
@@ Acknowledgements
This work was supported by JST SPRING, Grant Number JPMJSP2110 (YZ). This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, and JST CREST JPMJCR21N3 (HS). Additionally, we extend our gratitude to Kenji Fukumizu and Kei Hirose for engaging in insightful discussions and to the anonymous reviewers for their constructive feedback.
@@ A.1 More information about Experiment setup
WN18RR is a subset of WN18 Dettmers et al. (2018) which is contained in WordNet Miller (1995). FB15K-237 is a subset of FB15K, which is a subset of Freebase Bollacker et al. (2008), a comprehensive KG including data about common knowledge. All three datasets were designed for KGE, and we employ them for KGE tasks, and all three datasets have no individual people or offensive content.
@@ A.4 Introduction of Relation Patterns
Certain relations demonstrate symmetry, indicating that the validity of a relation between entities $x$ and $y$ ($\left(r_{1}(x,y)\Rightarrow r_{1}(y,x)\right)$) (for instance, is married to) is equally valid in the opposite direction (namely, from $y$ to $x$). Conversely, other relations are characterized by antisymmetry ($\left(r_{1}(x,y)\Rightarrow\neg r_{1}(y,x)\right)$), signifying that if a relation is applicable between $x$ and $y$ (such as is father of), it is inapplicable in the reverse direction (from $y$ to $x$).
@@
HittERChen et al. (2020) Transformer structure .503
@@ A.5 Other baseline KGE model
In recent times, several significant performance methods have been developed, as detailed for WN18RR in Table 6. Among these, MoCoSAHe et al. (2023), SimKGCWang et al. (2022a), C-LMKEWang et al. (2022b), and KNN-KGEZhang et al. (2022) mainly utilize Language Models (LMs) to enrich dataset semantic information, thereby achieving superior outcomes. Conversely, LERPHan et al. (2023) does not employ LMs but leverages additional contextual information (logic rules) beyond the dataset to fill information gaps in entities and relations, HittER Chen et al. (2020) integrates the transformer architecture into KGE, yet its lack of explainability remains unresolved. On the other hand, methods such as TransEBordes et al. (2013), RotatESun et al. (2019), and the OrthogonalE method introduced in this paper depend solely on the inherent data and information of the KGE dataset itself. These methods, based on specific mathematical rules and algorithms, do not incorporate any external information and thus do not operate as black-box approaches like LLMs. Consequently, these dataset-dependent methods remain highly valuable for KGE research.
@@
All the hyperparameter settings have been shown in Table 5.